Directrix (directrix):

How many squares are there on a checkerboard? (The answer is not 64.)

OpenStudy (anonymous):

1^2 +2^2 +3^2+.... + 8^2= 8*9*17/6=204

OpenStudy (anonymous):

@directrix, u know the answer?

OpenStudy (anonymous):

(n(n+1)(2n+1))/6 , let n = 8

Directrix (directrix):

@Arnabo09 Why do you ask?

OpenStudy (anonymous):

then tell me if its correct or not.. i guess u know the answer

Directrix (directrix):

I would like to get input from everyone who wants to reply. By then, I will have finished counting on my checkerboard. :)

OpenStudy (anonymous):

what i also have found is that a similar approach works for an (albeit imaginary) 3D "chess cube" the sum of the cubes: (n(n+1)/2)^2

OpenStudy (anonymous):

oh.. ok.. there will be 8 squares in a row if its length is 1 unit 7 squares in a row if its length is 2 unit.. 1 square in a row if its length is 8 unit.. so, total: 1^2+2^2+3^2+..+8^2

OpenStudy (anonymous):

what happens if we had a chessboard which isnt square? instead of NxN , we had NxM ?

OpenStudy (anonymous):

well, i guess if chessboard is mentioned, it is assumed its 8x8

OpenStudy (anonymous):

overlaping or not?

OpenStudy (anonymous):

well the problem is solved, i was just suggesting an extension to it

OpenStudy (anonymous):

if overlaping: 1+2*2+3*3+4*4+5*5+6*6+7*7+8*8

OpenStudy (anonymous):

then the numbers shouldnt be squared.. say, if its 8x6 1 unit length: 8*6 2 unit length:7*5 6 unit length: 3*1.. add these up, @eigen..

OpenStudy (anonymous):

if they are alowed to overlap each other, my answer is correct

OpenStudy (anonymous):

if you whant i explain....

OpenStudy (anonymous):

if not overlaping, you are right

Directrix (directrix):

204. There are 8 different sizes of squares, ranging from 1 × 1 to 8 × 8. Only 1 square is 8 × 8, but 4 squares are 7 × 7 and 9 squares are 6 × 6; this pattern continues through to 64 1 × 1 squares.

OpenStudy (callisto):

I think it is 1 + 4+9+16+25+36+49 +64 ..

OpenStudy (callisto):

Oh fail, I'm too late :S